evaluate is m=-2 and n=3: m^-2+n^-1

-2^-2+3^-1 = -8+(-3) = -11?

Answer by edjones(8007)

(-2)^(-2)+3^(-1)
=1/(-2)^2 + 1/3
=1/4 + 1/3
=3/12 + 4/12
=7/12
.
Ed